OpenStudy (anonymous):

Solve the proportion 47/39 =b/2

OpenStudy (anonymous):

(2)(47) = b(39)

OpenStudy (mayankdevnani):

\[\huge \frac{47}{39}=\frac{b}{2}\] do cross multiply, \[\huge 94=39b\] \[\huge b=\frac{39}{94}\]